# EasyGen: Easing Multimodal Generation with a Bidirectional Conditional Diffusion Model and LLMs

### v2

@@ Abstract
EasyGen: Easing Multimodal Generation with a Bidirectional Conditional Diffusion Model and LLMs
Xiangyu Zhao, Bo Liu [^1] Equal contribution., Qijiong Liu [^1] Equal contribution., Guangyuan Shi [^1] Equal contribution., Xiao-Ming Wu Department of Computing, The Hong Kong Polytechnic University {xiang-yu.zhao, bokelvin.liu, jyonn.liu, guang-yuan.shi}@connect.polyu.hk, xiao-ming.wu@polyu.edu.hk
@@ 4.2 Pre-training an Adapter to Enhance BiDiffuser’s SUR Capability
BiDiffuser uses the text encoder of CLIP, which is trained with image-text contrastive learning, limiting its semantic understanding and reasoning (SUR) ability for image generation. Drawing inspiration from Zhong et al. (2023), we utilize LLMs to enhance the SUR capability of LLMs. Specifically, we design an adapter that employs the attention mechanism to integrate the semantic information from LLM’s last hidden state $f_{\text{LLM}}(\cdot)$ into the CLIP text encoder $f_{\text{CLIP}}(\cdot)$. The adapter consists of a projection layer $\text{MLP}(\cdot)$ and a cross-attention layer $\text{Att}(\cdot)$. Given a paired image-text data $(\mathbf{x}_{0},\mathbf{y}_{0})$, we can get $y_{\text{sur}}$ with enhanced SUR via the adapter:
@@
$y_{0}=\lambda\cdot y_{\text{sur}}+(1-\lambda)\cdot f_{\text{CLIP}}(\mathbf{y}_{0}),$
@@ 4.4 Text-to-Image Response Generation
To tackle multimodal response generation tasks such as PhotoChat (Zang et al., 2021), we first finetune the LLM to generate detailed image captions based on dialogue context. Then, we employ BiDiffuser to create the corresponding images with the produced captions. Specifically, we replace the image featured in the dialogue with its corresponding descriptive caption, encapsulating it with task-specific tokens <Img>,</Img> and constructing the following instruction templates:
@@
where $w^{c}=(w_{1}^{c},...,w_{L}^{c})$ represents the caption of image $x_{0}$ with length $L$, $H$ is the dialogue history, and $\vartheta$ denotes the LLM’s parameters. Considering the potential for alignment deviation in text alone, we utilize the output of the LLM as the conditional component of the diffusion model. This is achieved through the adapter, which directly contributes to the denoising process. The loss function for the denoising process of a noisy image $x_{t^{x}}$ is formulated in a way that is similar to Eq. 4:
@@ 5.1 Experimental Setup
MLLM Sample Size Cosine Similarity $\uparrow$ MSE $\downarrow$ MiniGPT-4 5M 0.0016 6.2031 LLaVA v1.5 558K -0.0026 0.8433 Emu 2B 0.0054 0.4062 EasyGen 169K 0.0128 0.0338 Table 3: Data efficiency. Avg. Cosine similarity and mean square error between the projected representations and their respective captions embedded by LLM.
@@ 5.5 Data Efficiency in Training
MLLM FID $\downarrow$ Diffusion Model FID $\downarrow$ NExT-GPT 11.28 (+0.07) SD 11.21 Emu 11.66 (+1.73) SD v1.5 9.93 Fine-tuned on MS-COCO EasyGen 7.68 (-0.44) UniDiffuser 8.12 + w/o adapter 7.89 (-0.23) UniDiffuser 8.12 Fine-tuned on original dataset (zero-shot) EasyGen 9.16 (-0.55) UniDiffuser 9.71 + w/o adapter 9.52 (-0.19) UniDiffuser 9.71 Table 6: Comparing the image generation quality of end-to-end MLLMs and their corresponding diffusion models on the MS-COCO validation set (256 × 256). Our EasyGen surpasses the original diffusion model, while other MLLMs fall short in comparison.
@@ 5.6 Image Generation Quality
Tab. 6 evaluates the generated image’s quality of MLLMs on MS-COCO validation set, using 30K randomly selected prompts to compute the FID score on generated images. To confirm the efficacy of our approach, we fine-tuned our method on a portion of the original data (LIAON-COCO) and the MS-COCO train set, respectively. While other models resulted in a decrease in image generation performance compared to the corresponding diffusion model, EasyGen outperformed the original UniDiffuser due to the fine-tuned BiDiffuser and the adapter module. This suggests that the alignment between LLM and diffusion model’s text space used in other MLLMs may be unnecessary. Instead, generating images from texts/captions and aligning them with the diffusion model’s image space appears to be an effective solution.
@@ Appendix F Instruction Tuning
Different Instruction Templates: [Uncaptioned image] Caption Generation Provide a brief description of the given image. Assistant: Response Buses parked with a snow mountain view behind them. VQA What numbers are displayed on the front of the bus on the right? Short answer: Assistant: Response 6044 Multimodal Dialogue What might be the purpose of the buses in this location? Assistant: Response It is plausible that they are in this location for multiple reasons. Some possible reasons might be: 1)…2)… Table 11: Examples of different instructions with different output formats. We use bold fonts to indicate different instructions.

### v3

# EasyGen: Easing Multimodal Generation with BiDiffuser and LLMs

@@ Abstract
EasyGen: Easing Multimodal Generation with BiDiffuser and LLMs
Xiangyu Zhao, Bo Liu [^1] Co-Second Author., Qijiong Liu [^1] Co-Second Author., Guangyuan Shi [^1] Co-Second Author., Xiao-Ming Wu Department of Computing, The Hong Kong Polytechnic University {xiang-yu.zhao, bokelvin.liu, jyonn.liu, guang-yuan.shi}@connect.polyu.hk, xiao-ming.wu@polyu.edu.hk
@@ 4.2 Pre-training an Adapter to Enhance BiDiffuser’s SUR Capability
BiDiffuser uses the text encoder of CLIP, which is trained with image-text contrastive learning, limiting its semantic understanding and reasoning (SUR) ability for image generation. Drawing inspiration from Zhong et al. (2023), we utilize LLMs to enhance the SUR capability of BiDiffuser. Specifically, we design an adapter that employs the attention mechanism to integrate the semantic information from LLM’s last hidden state $f_{\text{LLM}}(\cdot)$ into the CLIP text encoder $f_{\text{CLIP}}(\cdot)$. The adapter consists of a projection layer $\text{MLP}(\cdot)$ and a cross-attention layer $\text{Att}(\cdot)$. Given a paired image-text data $(\mathbf{x}_{0},\mathbf{y}_{0})$, we can get $y_{\text{sur}}$ with enhanced SUR via the adapter:
@@
$y_{0}=\lambda\cdot y_{\text{sur}}+(1-\lambda)\cdot f_{\text{CLIP}}(\mathbf{y}_{0}),$ (4)
@@ 4.4 Text-to-Image Response Generation
To tackle multimodal response generation tasks such as PhotoChat (Zang et al., 2021), we first leverage the MLLM to generate detailed image captions based on dialogue context. Then, we employ BiDiffuser to create the corresponding images with the produced captions. Specifically, we replace the image featured in the dialogue with its corresponding descriptive caption, encapsulating it with task-specific tokens <Img>,</Img> and constructing the following instruction templates:
@@
where $w^{c}=(w_{1}^{c},...,w_{L}^{c})$ represents the caption of image $x_{0}$ with length $L$, $H$ is the dialogue history, and $\vartheta$ denotes the LLM’s parameters. Considering the potential for alignment deviation in discrete text alone, given the description of the image $\mathbf{y}_{0}$, we utilize $y_{0}$, which is the combination of the SUR adapter’s output and the CLIP text encoder’s output, as the conditional component of the diffusion model. This directly contributes to the denoising process. The loss function for the denoising process of a noisy image $x_{t^{x}}$ is formulated in a way that is similar to Eq. 5:
@@ 5.1 Experimental Setup
MLLM Sample Size Cosine Similarity $\uparrow$ MSE $\downarrow$
MiniGPT-4 5M 0.0016 6.2031
LLaVA v1.5 558K -0.0026 0.8433
Emu 2B 0.0054 0.4062
EasyGen 169K 0.0128 0.0338
Table 3: Data efficiency. Avg. Cosine similarity and mean square error between the projected representations and their respective captions embedded by LLM.
@@ 5.5 Data Efficiency in Training
MM-Model FID $\downarrow$ Diffusion Model FID $\downarrow$
Zero-Shot
NExT-GPT 11.28 (+0.07) SD 11.21
Emu 11.66 (+1.73) SD v1.5 9.93
EasyGen 9.16 (-0.55) UniDiffuser 9.71
+ w/o adapter 9.52 (-0.19) UniDiffuser 9.71
Fine-tuned on MS-COCO
EasyGen 7.68 (-0.44) UniDiffuser 8.12
+ w/o adapter 7.89 (-0.23) UniDiffuser 8.12
Table 6: Comparing the image generation quality of end-to-end MLLMs and their corresponding diffusion models on the MS-COCO validation set (256 × 256). Our EasyGen surpasses the original diffusion model, while other MLLMs fall short in comparison.
@@ 5.6 Image Generation Quality
Tab. 6 evaluates the generated image’s quality of MLLMs on MS-COCO validation set, using 30K randomly selected prompts to compute the FID score on generated images. To confirm the efficacy of our approach, we fine-tuned our method on a portion of the original data (LIAON-COCO) and the MS-COCO train set, respectively. While other models resulted in a decrease in image generation performance compared to the corresponding diffusion model, EasyGen outperformed the original UniDiffuser due to the fine-tuned BiDiffuser and the adapter module. Furthermore, Tab. 7 provides CLIP-T scores from ImagenHub. We notice similar trends to the results in Tab. 6 using the FID indicator. This suggests that our method can better align LLM with diffusion model’s text space.
MM-Model CLIP-T $\uparrow$ Diffusion Model CLIP-T $\uparrow$
NExT-GPT 0.259 (-0.031) SD 0.290
Emu 0.262 (-0.023) SD v1.5 0.285
Emu2 0.266 (-0.023) SD XL 0.289
EasyGen 9.16 (-0.55) UniDiffuser 9.71
Table 7: Comparing the CLIP-T score of end-to-end MLLMs and their corresponding diffusion models on the ImagenHub.
@@ Appendix F Instruction Tuning
Different Instruction Templates:
Caption Generation Provide a brief description of the given image. Assistant:
Response Buses parked with a snow mountain view behind them.
VQA What numbers are displayed on the front of the bus on the right? Short answer: Assistant:
Response 6044
Multimodal Dialogue What might be the purpose of the buses in this location? Assistant:
Response It is plausible that they are in this location for multiple reasons. Some possible reasons might be: 1)…2)…
Table 12: Examples of different instructions with different output formats. We use bold fonts to indicate different instructions.